# Retrieval Pivot Attacks in Hybrid RAG: Measuring and Mitigating Amplified Leakage from Vector Seeds to Graph Expansion

### v1

@@ XI-A Vector RAG Poisoning
PoisonedRAG [28] formalized knowledge corruption attacks on RAG, achieving 90% ASR with 5 injected documents. CorruptRAG [24] demonstrated single-document attacks with higher stealth. CtrlRAG [18] achieved 90% ASR on GPT-4o via black-box feedback optimization. RIPRAG [21] applied reinforcement learning to optimize poisoning without model access. NeuroGenPoisoning [27] targeted specific neurons for ¿90% success. These works study vector-side attacks in isolation; none address what happens when poisoned vector results seed graph expansion.
XI-B GraphRAG Security
@@
GRAGPoison [8] is the closest related work, demonstrating relation-centric poisoning of GraphRAG with 98% ASR. TKPA [20] showed that modifying 0.06% of corpus text drops GraphRAG accuracy by 45 percentage points. RAGCrawler [22] achieved 84.4% knowledge extraction coverage through graph-guided probing. The Graph RAG Privacy Paradox [10] established that graph RAG increases structural leakage while reducing text leakage. Our work extends these graph-side insights to the hybrid setting, where the vector-to-graph transition creates a compound attack surface.
XI-C RAG Privacy and Extraction
@@
The SoK on RAG Privacy [2] systematized all known privacy attack vectors in RAG systems and explicitly noted that hybrid RAG privacy risks remain under-studied. Riddle Me This [12] demonstrated membership inference on RAG systems. Traceback of RAG Poisoning [23] provided forensic methods for identifying responsible documents, but traces only through the vector retrieval path without following graph expansion chains.
XI-D RAG Defenses
@@
RAGuard [3] detects poisoning via retrieval pattern analysis. SeCon-RAG [17] applies semantic consistency filtering. RevPRAG [19] achieves 98% detection via LLM activation analysis. SDAG [5] partitions context into trusted and untrusted segments with block-sparse attention. SD-RAG [1] implements selective disclosure policies. All operate within a single retrieval modality. Our defense suite (D1–D5) is the first designed specifically for the cross-store boundary, with per-hop authorization as the cornerstone mechanism.
XI-E Ethical Considerations
@@
Experiments use two corpora: a synthetic corpus generated by the authors and the Enron email corpus, which is a public-record dataset released during the 2001 FERC investigation and widely used in NLP research [4]. No production enterprise data, user accounts, or confidential documents were used. The attack implementations (A1–A7) target our own evaluation infrastructure and are designed to characterize vulnerabilities, not exploit production systems. Our primary contribution is the defense (D1), which we release alongside the attack code to ensure that the net effect of publication is protective. The released code does not include tools for targeting external systems; all components require a local Neo4j and ChromaDB deployment to operate.
XII Conclusion
@@
Hybrid RAG pipelines that combine vector retrieval with knowledge graph expansion introduce an attack surface at the vector-to-graph boundary that has received limited explicit treatment in prior work. We formalized this threat as Retrieval Pivot Risk (RPR) and demonstrated across two corpora—a synthetic enterprise dataset (1,000 documents, 4 tenants) and the Enron email corpus (50,000 emails, 5 departments)—that undefended hybrid pipelines exhibit $\text{RPR}=0.95$ and $\text{RPR}=0.70$ respectively, compared to $0.0$ for vector-only retrieval. All leakage occurs at 2 hops through the entity-linking pivot in our bipartite graph construction, a structural invariant that holds across both corpora.
@@
Our taxonomy of seven Retrieval Pivot Attacks (A1–A4 non-adaptive, A5–A7 adaptive) shows that small injection budgets (10–20 chunks) can exploit this boundary, and that even without adversarial injection, naturally shared entities create organic pivot paths that leak cross-tenant data. The Enron corpus confirms this on real data: 19 naturally shared entities across departmental boundaries produce RPR == 0.695 on benign queries alone.
@@
The most important finding is practical: per-hop authorization—re-checking tenant and sensitivity labels at each graph expansion step—eliminates all measured retrieval pivot leakage across both corpora, all seven attack variants, and metadata forgery rates up to 10%. This defense is simple, requires no model changes, adds <<1ms latency, and uses metadata already present in graph databases. We recommend it as the minimum viable security control for any hybrid RAG deployment.

### v2

@@ 11.1 Vector RAG Poisoning
PoisonedRAG [zou2025poisonedrag] formalized knowledge corruption attacks on RAG, achieving 90% ASR with 5 injected documents. CorruptRAG [corruptrag2025] demonstrated single-document attacks with higher stealth. CtrlRAG [ctrlrag2025] achieved 90% ASR on GPT-4o via black-box feedback optimization. RIPRAG [riprag2025] applied reinforcement learning to optimize poisoning without model access. NeuroGenPoisoning [neurogenpoison2025] targeted specific neurons for >90% success. These works study vector-side attacks in isolation; none address what happens when poisoned vector results seed graph expansion.
11.2 GraphRAG Security
@@
GRAGPoison [gragpoison2026] is the closest related work, demonstrating relation-centric poisoning of GraphRAG with 98% ASR. TKPA [fewwords2025] showed that modifying 0.06% of corpus text drops GraphRAG accuracy by 45 percentage points. RAGCrawler [ragcrawler2026] achieved 84.4% knowledge extraction coverage through graph-guided probing. The Graph RAG Privacy Paradox [graphragprivacy2025] established that graph RAG increases structural leakage while reducing text leakage. Our work extends these graph-side insights to the hybrid setting, where the vector-to-graph transition creates a compound attack surface.
11.3 RAG Privacy and Extraction
@@
The SoK on RAG Privacy [sokprivacy2026] systematized all known privacy attack vectors in RAG systems and explicitly noted that hybrid RAG privacy risks remain under-studied. Riddle Me This [riddleme2025] demonstrated membership inference on RAG systems. Traceback of RAG Poisoning [traceback2025] provided forensic methods for identifying responsible documents, but traces only through the vector retrieval path without following graph expansion chains.
11.4 RAG Defenses
@@
RAGuard [raguard2025] detects poisoning via retrieval pattern analysis. SeCon-RAG [seconrag2025] applies semantic consistency filtering. RevPRAG [revprag2025] achieves 98% detection via LLM activation analysis. SDAG [sdag2026] partitions context into trusted and untrusted segments with block-sparse attention. SD-RAG [sdrag2026] implements selective disclosure policies. All operate within a single retrieval modality. Our defense suite (D1–D5) is the first designed specifically for the cross-store boundary, with per-hop authorization as the cornerstone mechanism.
Ethical Considerations
@@
Experiments use three corpora: a synthetic corpus generated by the authors, the Enron email corpus (a public-record dataset released during the 2001 FERC investigation and widely used in NLP research [enroncorpus2015]), and SEC EDGAR 10-K filings (publicly available regulatory documents). No production enterprise data, user accounts, or confidential documents were used. The attack implementations (A1–A7) target our own evaluation infrastructure and are designed to characterize vulnerabilities, not exploit production systems. Our primary contribution is the defense (D1), which we release alongside the attack code to ensure that the net effect of publication is protective. The released code does not include tools for targeting external systems; all components require a local Neo4j and ChromaDB deployment to operate.
Open Science
We release the complete research artifact: source code, synthetic data generators, Enron ingestion pipeline, all seven attack implementations (A1–A7), the five-layer defense suite (D1–D5), evaluation harness, query templates, and raw experimental results. The repository includes 255 passing unit tests and deterministic reproduction via a fixed random seed (42). All experiments run on commodity hardware (<<16 GB RAM, CPU-only) using open-source models and databases (spaCy, Sentence Transformers, Neo4j Community Edition, ChromaDB). The Enron email corpus is a public-record dataset available from Carnegie Mellon University [enroncorpus2015]. No proprietary models, commercial APIs, or restricted datasets are required to reproduce the core retrieval pivot results. The generation contamination evaluation (§7.8) uses commercial LLM APIs (OpenAI, Anthropic, DeepSeek) and is therefore dependent on API availability and pricing; we include cached results for reproducibility without API access.
12 Conclusion
@@
Hybrid RAG pipelines that combine vector retrieval with knowledge graph expansion introduce an attack surface at the vector-to-graph boundary that has received limited explicit treatment in prior work. We formalized this threat as Retrieval Pivot Risk (RPR) and demonstrated across three corpora—a synthetic enterprise dataset (1,000 documents, 4 tenants), the Enron email corpus (50,000 emails, 5 departments), and SEC EDGAR 10-K filings (887 sections, 4 sector-tenants)—that undefended hybrid pipelines exhibit $\text{RPR}=0.95$, $0.70$, and $0.09$ respectively, compared to $0.0$ for vector-only retrieval. All leakage occurs at 2 hops through the entity-linking pivot in our bipartite graph construction, a structural invariant that holds across all three corpora.
@@
Our taxonomy of seven Retrieval Pivot Attacks (A1–A4 non-adaptive, A5–A7 adaptive) shows that small injection budgets (10–20 chunks) can exploit this boundary, and that even without adversarial injection, naturally shared entities create organic pivot paths that leak cross-tenant data. The Enron and EDGAR corpora confirm this on real data: 19 naturally shared entities across Enron departmental boundaries produce RPR == 0.695, while 13 cross-sector entities in EDGAR 10-K filings produce RPR == 0.085 on benign queries alone.
@@
The most important finding is practical: per-hop authorization—re-checking tenant and sensitivity labels at each graph expansion step—eliminates all measured retrieval pivot leakage across all three corpora, all seven attack variants, and metadata forgery rates up to 10%. This defense is simple, requires no model changes, adds <<1ms latency, and uses metadata already present in graph databases. We recommend it as the minimum viable security control for any hybrid RAG deployment.